# Fundamental mechanisms of high-electric-field vacuum arc initiation

### v1

@@ II.2 Radio Frequency experiments
The experimental equipment, operating conditions and procedures used to determine the high-gradient performance and long-term operating behavior of radio frequency accelerating structures as well as experiments dedicated to the fundamental understanding of high-field phenoma are described in this section. The key experimental measureables including field emission, breakdown, conditioning and breakdown rate and how they are measured will be introduced.
@@
The radio frequency experiments and experimental infrastructure are described in some detail in order to provide a solid basis for understanding the data used for developing the concepts described in this report. In addition, the description of the radio frequency system is important as an introduction to section IX which describes the critical role the coupling of power has on the evolution and, it turns out, the onset level of breakdown. The result is an important variation of maximum achievable surface fields as a function of structure geometry in the case of a radio frequency structure, and powering circuits in the case of dc systems.
@@
A short description will now be made of the basic operating principle, geometry and electromagnetic field pattern of an RF accelerating structure. This is can be helpful for many readers since radio frequency structures do not have clearly separated anodes and cathodes as there are in the dc high-field systems systems that have been the focus of most breakdown studies. The properties of radio frequency structures are described in more depth in section IX.
@@
The basic configuration of an accelerating radio frequency structure is a set of longitudinally coupled pillbox-like cavities, operating in the TM010 mode. The coupling irises also allow for passage of the beam. A schematic examples of the geometry and electric field in accelerating cells are shown in Fig. 1. The longitudinal electron field, the middle image in Fig. 1 provides the accelerating force on a bunched charged particle beam when the beam arrives with the correct phase. When coupled together with the correct phase relationship, a synchronous acceleration is obtained over multiple cavities. Multiple cavities can be coupled either in a so-called standing wave configuration, in which there is a single input power feed, and traveling wave, in which there is both an input and an output power feed. In a standing wave structure the field builds up over the entire structure with time after the beginning of a power pulse while in a traveling wave structure the power pulse travels along the length of the structure.
@@
The phase relationship between successive cells, the phase advance, must be matched to the velocity of the particle beam in order to provide synchronous acceleration over multiple cells. The phase advance in in both standing wave and traveling wave structures is determined by the frequency of the cavities, which is primarily a function of the diameter of the cavity, and the coupling strength between the successive cells, determined primarily by the diameter and thickness of the coupling iris. The coupling between the cells also determines the mode separation in standing wave structures and the group velocity in traveling structures.
@@
The geometry of an RF accelerating structure is consequently strongly influenced by the requirements given by synchronous acceleration. Further influence on the geometry are given by the efficiency and power-effectiveness of the acceleration. As we will see these acceleration-driven influences often go in opposition to those needed for maximizing field-holding. A fundamental understanding of high-field phenomena can have a significant role in resolving this optimization. We will also see that the field and power coupling that determines the power flow characteristics in the structures also similarly determines the coupling of the power to an incipient breakdown, and through this determines the field holding capability of a structure. This idea will be elaborated in depth in section IX.
@@
The surface electric field in an accelerating structure is concentrated on the inner edge of the coupling iris as shown in the lower image in Fig. 1. This is because the flux lines that would normally fall onto the center area of a pillbox cavity operating in the TM010 mode cannot due to the absence of metal, so they connect to the closest cavity wall which is the coupling iris. Because of this electric field concentration, most vacuum breakdowns are concentrated on the inner edges of the coupling irises.
@@
The operating frequency of linear accelerators and test stands typically lies in the S, C, and X-band, along with Ka band during the first stage of the CLIC project. The electromagnetic fields inside an accelerating structure are established by feeding it with radio frequency power via a transmission line and through an impedance-matching coupler. The transmission line is typically a hollow rectangular waveguide in high-frequency systems.
In high-gradient experiments and linacs in general, the RF system operates in a pulsed mode. In the case of the CLIC linear collider the radio frequency pulses going into the accelerating structures have a power and pulse length of around 50 MW and 200 ns respectively, for a pulse energy of around 10 J, and are produced at repetition rates in the range of 50 to 100 Hz. This report focuses on vacuum breakdown in pulsed operation although connections to systems with continuous field exposure are discussed in the section III.
The electric and magnetic field levels inside an radio frequency device depend on the geometry of the device and are proportional to the square root of the input power. These radio frequency properties can be determined using commercial electromagnetic solvers such as HFSS and CST Microwave studio. The discussion in this section will focus on traveling wave structures since they have made up the bulk of the test devices that underlie analysis described in this review. Power flow considerations in the case of breakdowns in standing wave structures is taken up in the breakdown power coupling section IX.
Vacuum breakdowns affect the transmission through, and reflection from the accelerating structure. In particular breakdowns cause abrupt, ns-scale, changes in power flow with suppression of transmission and large reflected signals. We will return to this point in Subsec. III.2. Consequently power flows are one of the main experimental signals. They are sampled through directional couplers or similar devices and are monitored using detector systems that typically consist of power-sensitive diodes, or heterodyne receivers, and signal digitization. The acquisition is typically done pulse by pulse, with a sequence of pulses stored in a buffer.
@@
One of the most important measures of the operation of an accelerating structure is its breakdown rate. The breakdown rate is defined as the number of pulses with breakdown divided by the total number of pulses in measurement window. The breakdown rate is strongly dependent on field level, as will be described in depth in section III.3. Pulsing is typically inhibited after a pulse with a breakdown has been detected. A short time, typically of the order of a second, is given for vacuum levels to return to normal and pulsing starts again, initially at a lower power level which is gradually returned to nominal.
When a breakdown occurs a large current is emitted from the breakdown site. As we will see in the breakdown power coupling section IX this current accelerated in the radio frequency structure fields is what absorbs incident power and causes reflected power. Most of the breakdown current is accelerated into the radio frequency structure walls, however some of it is accelerated with a longitudinal component and exits the input and output end of the structure. The measurement of this breakdown current is detected using Faraday cups that are placed outside an accelerating structures under test along the beam axis. On-axis Faraday cups can only be implemented in test stands that do not accelerate an externally injected beam.
@@
Pulsed dc experiments have been carried out to complement radio frequency experiments with a system that is simpler, less expensive and easier to instrument. In a DC experiment voltage is applied across two clearly defined electrodes: cathode and anode. This is an obvious advantage when investigating breakdown compared to RF experiments where the dynamics have a further degree of freedom due to the fast polarity changes. In addition the electrode geometry can simpler than an RF one so localizing breakdown is simpler.
@@
The first dc setup of the CLIC study consisted of a tip facing a plane sample, with the former as the anode and the latter as the cathode. In such a configuration, and with a tip of the order of the size of a breakdown plasma spot, the breakdown location is defined by the tip shape and dimensions. The experiments performed at CERN were performed with an anode tip with an apex radius of the order of separated from the sample by a gap of , accurately regulated with a micro-manipulator (Fig. 4). The tip and sample were placed in an all-metal ultra-high vacuum (UHV) system, pumped with turbo-molecular pumps down to less than and instrumented with standard vacuum gauges and a residual gas analyser [77]. This type of experiments draws inspiration from the large body of previous literature, in particular for the choice of having the test sample setup as the cathode. Evidence in the literature, as in this review, is that breakdowns start on the cathode and that critical aspects such as the threshold field depends solely on the cathode material. This will be discussed extensively later in this review.
@@
A typical setup of the anode tip system consisted of a capacitor bank charged by a dc power supply, as shown in Fig. 5. Once the capacitor bank was charged, it was connected either by mechanical or electronic switches to the electrode allowing for a fast rise of the potential (often in the range), accompanied by a displacement-current surge due to the intrinsic and stray capacitance of the system and its connecting cables. The high voltage was held for a defined time interval, typically from a few hundreds of to hundreds of , or even longer intervals up to seconds, depending on the type of experiment. When a breakdown happens it manifests itself through a discharge current of several tens or hundreds of , which is easily detected by the control electronics and found typically to be in coincidence with pressure spikes in the vacuum system [33] and with optical emission from the plasma [82]. In case of an arc the capacitor bank is entirely discharged through it; in case an arc did not develop during the programmed pulse duration, the capacitor bank is discharged on a dumping resistance. The cycle was then restarted, either at the same field, higher, or lower, depending on the type of experiment. The arc current depended mostly on the energy stored on the capacitor bank (which in turn depends also on the applied voltage) and on a possible matching series resistance [163], knowing that voltage across a vacuum arc and thus its electrical resistance has a typical dependence on the cathode material linked to its first ionization potential, of the order of for copper [5]. Repetition rate of such cycles ranged from several seconds per cycle in the earliest experiments, to several when more performing solid-state electronics was adopted.
@@
An important result was the demonstration of the similarity observed by scanning electron microscopy (SEM) [76, 163] of the breakdown craters on the cathode of the pulsed DC system and on the surface of RF cavities for a similar stored energy, as will be discussed later in III.8. This confirmed the relevance of these experiments to the study of high-gradient RF accelerating structures.
@@
Another important experiment that was routinely performed is measurement of field emission current. In this case, the high-voltage power supply is connected directly to the system via a very large limiting resistor, in the order of several tens of , commensurate with the expected resistance of the gap. Due to the extremely low measured currents, in the order of up to a few , it is important to prevent the occurrence of a high-current breakdown that could destroy the electronics making use of to such limiting resistor. From the field emission it is then possible to extract standard theoretical parameters such as the Fowler-Nordheim field enhancement factor . This measurement is performed at various stages of the experiments, and in some cases even before and after each high-voltage pulse in order to identify the evolution of at each pulse and breakdown [33]. Field enhancement factor for various metals has been measured, and ranges typically between 30 and 50 for different metals and at different conditioning stages [76, 34]. Field emission data is presented in section III.7.
@@
Moving from mechanical switches to fast solid state electronics in the subsequent generations of the test setup allowed pulse repetition rates increase from sub-Hz to several Hz and later to several kHz. At the same time, the micron-accuracy manufacturing technologies devised for the RF structures had become mature enough to allow fabrication of large components at moderate cost. As a consequence, a test system where two large flat electrodes was constructed. The electrodes are kept separated by a gap of a few tens of microns by an insulating ceramic spacer. The electrodes are exposed to high-repetition DC high-field pulses for exploring the breakdown rate behaviour of large surfaces, adding the statistic and stochastic nature of breakdown development to the scope of the study. This Large Electrode System (LES), also known as the pulsed dc system, is described in detail in [143], and is illustrated in Fig. 6.
The pulsed dc system consists of a UHV chamber where two large electrodes (of various diameters depending on the experiments, up to of active surface) are hosted, separated by a precision machined ceramic spacer maintaining a gap between the two active surface (typically , although several spacers are available allowing gaps ranging from to ). The most recent version of this test system is supplied by a Marx pulse generator, specially developed for the purpose [147]. Depending on the experiment, either electrode can be connected either to positive or negative polarity, and pulses of arbitrary duration from a few 100s of ns to several 100s of can be generated with a rise-time of the order of , as illustrated in Fig. 7, where typical current and voltage traces are shown in the two cases when a breakdown is observed or not. The availability of very fast electronics with pulse repetition rate up to allowed working in very low breakdown rate regime ( or lower) and reproducing the conditioning process as performed in typical RF tests. In particular, a conditioning algorithm similar to what developed for RF X-band tests [184] could be developed and adapted to pulsed DC tests. Moreover, the system has been designed with viewports on the sides, which allowed the installation of CCD cameras to detect the light emitted by breakdowns, and to reconstruct their position through triangulation, as shown in Fig. 8. The location of breakdowns can then be easily correlated with their counterparts acquired through microscopic imaging of the electrodes surfaces, as illustrated in Fig. 9, and with the electrical signals acquired by the pulse generator to analyze for example their size and their spatial and temporal distributions. These diagnostics are complemented by the usual vacuum pressure recordings discussed for the tip-to-sample original tests system.
@@ III.2 Signals produced during breakdown
In this subsection we present data that illustrate features that are characteristic of breakdowns taken individually. These features include how applied fields collapse, the contemporaneous burst-like signals of electron, radiation and light emission, as well as vacuum spikes. The data shown are by necessity selected examples since most of the breakdown features are random and noisy. The examples are chosen to illustrate features clearly but with care so as not to bias the discussion. Some categories of randomness in the features has been identified to be important and these will be subsequently covered in the statistics subsection III.3.
@@
From the length of the drop-time in transmission and rise-time in reflection, as shown in Figs. 3 and 7, one can determine that the development time of a breakdown short circuit is very short, less than 10 ns from all available data. The time is given as an upper limit since the measurement is bandwidth limited by the structure, detection electronics and the signal acquisition.
@@
The electron current emitted during breakdown, the one causing the field collapse described above, can also be observed directly in RF accelerating structure experiments by Faraday cups installed along the beam axis. Most of the breakdown current is accelerated and then lost on the cavity walls, but a fraction is captured by the accelerating RF and is transported along the central beam axis, the beam aperture, out of the structure. It can then be directly measured using on-axis Faraday cups and connected to an oscilloscope. Example of such Faraday cup signals on normal pulses and pulses without are shown in [131].

### v2

@@ II.2 Radio Frequency experiments
The experimental equipment, operating conditions, and procedures used to determine the high-gradient performance and long-term operating behavior of radio frequency accelerating structures, as well as experiments dedicated to the fundamental understanding of high-field phenomena, are described in this section. The radio frequency infrastructure was designed to allow testing of prototype linear collider components, in particular accelerating structures, at full operating parameters. As testing proceeded, key behaviors, dependencies and experimental measureables including field emission, breakdown, conditioning and breakdown rate were identified. These effects will be described in the next section, III.
@@
The radio frequency experiments and experimental infrastructure are described in some detail in order to provide a solid basis for understanding the data used for developing the concepts described in this report. In addition, the description of the radio frequency system is important as an introduction to section IX, which describes the critical role the coupling of power has on the evolution and, it turns out, the onset level of breakdown. The result is an important variation of maximum achievable surface fields as a function of structure geometry in the case of a radio frequency structure, and powering circuits in the case of dc systems.
@@
A short description will now be made of the basic operating principle, geometry and electromagnetic field pattern of an rf accelerating structure. This can be helpful for many readers since radio frequency structures do not have clearly separated anodes and cathodes as there are in the dc high-field systems that have been the focus of most breakdown studies. The properties of radio frequency structures are described in more depth in section IX.
@@
The basic configuration of an accelerating radio frequency structure is a set of longitudinally coupled pillbox-like cavities, operating in the TM010 mode. The coupling irises also allow for passage of the beam. A schematic examples of the geometry and electric field in accelerating cells are shown in Fig. 2. The electron field, the middle image in Fig. 2 provides the accelerating force on a bunched charged particle beam when the beam arrives with the correct phase. When coupled together with the correct phase relationship, a synchronous acceleration is obtained over multiple cavities. Multiple cavities can be coupled either in a so-called standing wave configuration, in which there is a single input power feed, and traveling wave, in which there is both an input and an output power feed. In a standing wave structure, the field builds up over the entire structure with time after the beginning of a power pulse, while in a traveling wave structure, the power pulse travels along the length of the structure.
@@
The phase relationship between successive cells, the phase advance, must be matched to the velocity of the particle beam in order to provide synchronous acceleration over multiple cells. The phase advance in both standing wave and traveling wave structures is determined by the frequency of the cavities, which is primarily a function of the diameter of the cavity, and the coupling strength between the successive cells, determined primarily by the diameter and thickness of the coupling iris. The coupling between the cells also determines the mode separation in standing wave structures and the group velocity in traveling wave structures.
@@
The geometry of an rf accelerating structure is consequently strongly influenced by the requirements given by synchronous acceleration. Further influence on the geometry is given by the efficiency and power-effectiveness of the acceleration. A fundamental understanding of high-field phenomena can have a significant role in resolving this optimization. We will also see that the field and power coupling that determines the power flow characteristics in the structures also similarly determines the coupling of the power to an incipient breakdown, and through this determines the field-holding capability of a structure. This idea will be elaborated in depth in section IX.
@@
The surface electric field in an accelerating structure is concentrated on the inner edge of the coupling iris as shown in the lower image in Fig. 2. Because of this electric field concentration, most vacuum breakdowns are concentrated on the inner edges of the coupling irises.
@@
The operating frequency of linear accelerators and test stands typically lies in the S, C, and X-bands, along with Ka-band during the first stage of the CLIC project. The electromagnetic fields inside an accelerating structure are established by feeding it with radio frequency power via a transmission line and through an impedance-matching coupler. The transmission line is typically a hollow rectangular waveguide in high-frequency systems.
In high-gradient experiments and linacs in general, the rf system operates in a pulsed mode. In the case of the CLIC linear collider the radio frequency pulses going into the accelerating structures have a power and pulse length of around 50 MW and 200 ns respectively, for a pulse energy of around 10 J, and are produced at repetition rates in the range of 50 to 100 Hz. This report focuses on vacuum breakdown in pulsed operation, although connections to systems with continuous field exposure are discussed in the section III.
The electric and magnetic field levels inside a radio frequency device depend on the geometry of the device and are proportional to the square root of the input power. These radio frequency properties can be determined using commercial electromagnetic solvers such as HFSS and CST Microwave studio. The discussion in this section will focus on traveling wave structures since they have made up the bulk of the test devices that underlie analysis described in this review. Power flow considerations in the case of breakdowns in standing wave structures is taken up in the breakdown power coupling section IX.
Vacuum breakdowns affect the transmission through and reflection from the accelerating structure. In particular, breakdowns cause abrupt, nanosecond-scale changes in power flow with suppression of transmission and large reflected signals. We will return to this point in subsection III.2. Consequently, power flows are one of the main experimental signals. They are sampled through directional couplers or similar devices and are monitored using detector systems that typically consist of power-sensitive diodes, or heterodyne receivers, and signal digitization. The acquisition generally is done pulse by pulse, with a sequence of pulses stored in a buffer.
@@
One of the most important measures of the operation of an accelerating structure is its breakdown rate. The breakdown rate is defined as the number of pulses with breakdown divided by the total number of pulses in the measurement window. The breakdown rate is strongly dependent on field level, as will be described in depth in section III.3. Pulsing is typically inhibited after a pulse with a breakdown has been detected. A short time, typically of the order of a second, is given for vacuum levels to return to normal and pulsing starts again, initially at a lower power level which is gradually returned to nominal.
When a breakdown occurs a large current is emitted from the breakdown site. As we will see in the breakdown power coupling section IX, this current accelerated in the radio frequency structure fields is what absorbs incident power and causes reflected power. Most of the breakdown current is accelerated into the radio frequency structure walls, however some of it is accelerated with a longitudinal component and exits the input and output ends of the structure. The measurement of this breakdown current is detected using Faraday cups that are placed outside an accelerating structures under test along the beam axis. On-axis Faraday cups can only be implemented in test stands that do not accelerate an externally injected beam.
@@
The radio frequency devices described in this review typically operate in the range to mbar. The vacuum level must be continuously monitored since it tends to be high during initial operation, as described in subsection III.4 and pressure bursts are caused by breakdowns, as described in Subsec. III.2.
@@
Pulsed dc experiments have been carried out to complement radio frequency experiments using systems that are simpler, less expensive and easier to instrument. In addition pulsed dc experiments can produce results more quickly because electrodes are simpler to fabricate than a radio frequency structure and in later versions, were carried out at a much higher repetition rate. Overall the parameters of the pulsed dc experiments were chosen to overlap with the radio frequency in important aspects for example electrode material, heat treatment and machining and pulse length but also to extend beyond to provide a broader parameter space for benchmarking the theoretical work. Examples of the latter include pulse length, repetition rate and electrode temperature.
@@
In a dc experiment voltage is applied across two clearly defined electrodes: cathode and anode. This is an obvious advantage when investigating breakdown compared to rf experiments where the dynamics have a further degree of freedom due to the fast polarity changes. In addition, the electrode geometry can be simpler than an rf cavity so localizing breakdown is simpler.
@@
The first dc setup of the CLIC study consisted of a tip facing a plane sample, with the former as the anode and the latter as the cathode. In such a configuration, and with a tip of the order of the size of a breakdown plasma spot, the breakdown location is defined by the tip shape and dimensions. The experiments performed at CERN were performed with an anode tip with an apex radius of the order of separated from the sample by a gap of , accurately regulated with a micro-manipulator (Fig. 5). The tip and sample were placed in an all-metal ultra-high vacuum (UHV) system, pumped with turbo-molecular pumps down to less than and instrumented with standard vacuum gauges and a residual gas analyser [81]. This type of experiment draws inspiration from the large body of previous literature, in particular for the choice of having the test sample setup as the cathode. Evidence in the literature, as in this review, is that breakdowns start on the cathode and that critical aspects such as the threshold field depends solely on the cathode material. This will be discussed extensively later in this review.
@@
A typical setup of the anode tip system consisted of a capacitor bank charged by a dc power supply, as shown in Fig. 6 (represented in the lower branch). Once the capacitor bank was charged, it was connected either by mechanical or electronic switches to the electrode allowing for a fast rise of the potential (often in the range), accompanied by a displacement-current surge due to the intrinsic and stray capacitance of the system and its connecting cables. The high voltage was held for a defined time interval, typically from a few hundreds of to hundreds of , or even longer intervals up to seconds, depending on the type of experiment, and on the circuit hardware. When a breakdown happens, it manifests itself through a discharge current of several tens or hundreds of , which is easily detected by the control electronics and found typically to be in coincidence with pressure spikes in the vacuum system [36] and with optical emission from the plasma [87]. In this case the capacitor bank is entirely discharged through the arc, and the experimental cycle restarted by charging again the capacitor bank typically at a lower voltage. In case an arc did not develop during the programmed pulse duration, the capacitor bank is discharged on a dumping resistance. The cycle is then restarted, either at the same field in case of breakdown rate experiments, or higher field in case of ramp-to-breakdown experiments. The arc current depends mostly on the energy stored on the capacitor bank (which in turn depends also on the applied voltage) and on a possible matching series resistance [173]. The voltage across a vacuum arc, and thus its electrical resistance depends on the cathode material and is of the order of for copper [5]. Repetition rate of such cycles ranged from several seconds per cycle in the earliest experiments, to several when more performing solid-state electronics was adopted.
@@
An important result was the demonstration of the similarity observed by scanning electron microscopy (SEM) [80, 173] of the breakdown craters on the cathode of the pulsed dc system and on the surface of rf cavities for a similar stored energy, as will be discussed later in subsection III.8. This confirmed the relevance of these experiments to the study of high-gradient rf accelerating structures.
@@
Another important experiment that was routinely performed is the measurement of field emission current. In this case, the power supply is connected directly to the system via a limiting resistor and the FE current is read with a multimeter (represented in the top circuit branch of Fig. 6). The voltage is increased from zero to the value producing a FE current of , which is a safe value. With such a low FE current, a breakdown is not triggered and damaging the multimeter is avoided. From the field emission it is then possible to extract standard theoretical parameters such as the Fowler-Nordheim field enhancement factor . This measurement is performed at various stages of the experiments, and in some cases even before and after each high-voltage pulse in order to identify the evolution of at each pulse and breakdown [36]. Field enhancement factor for various metals has been measured, and ranges typically between 30 and 50 for different metals and at different conditioning stages [80, 37]. Field emission data is presented in subsection III.7.
@@
Moving from mechanical switches to fast solid state electronics in the subsequent generations of the test setup allowed pulse repetition rates increase from sub-Hz to several Hz and later to several kHz. At the same time, the micron-accuracy manufacturing technologies devised for the rf structures had become mature enough to allow fabrication of large components at moderate cost. As a consequence, a test system where two large flat electrodes were constructed. The electrodes are kept separated by a gap of a few tens of microns by an insulating ceramic spacer. The electrodes are exposed to high-repetition dc high-field pulses for exploring the breakdown rate behaviour of large surfaces, adding the statistical and stochastic nature of breakdown development to the scope of the study. This Large Electrode System (LES), also known as the pulsed dc system, has been described in detail in [152], and is illustrated in Fig. 7.
The pulsed dc system consists of a UHV chamber where two large electrodes (of various diameters depending on the experiments, up to of active surface) are hosted, separated by a precision machined ceramic spacer maintaining a gap between the two active surface (typically , although several spacers are available allowing gaps ranging from to ). The most recent version of this test system is supplied by a Marx pulse generator, specially developed for the purpose [156]. Depending on the experiment, either electrode can be connected either to positive or negative polarity, and pulses of arbitrary duration from a few 100s of ns to several 100s of can be generated with a rise-time of the order of , as illustrated in Fig. 8, where typical current and voltage traces are shown in the two cases when a breakdown is observed or not. The availability of very fast electronics with pulse repetition rate up to allowed working in very low breakdown rate regime ( or lower) and reproducing the conditioning process as performed in typical rf tests. In particular, a conditioning algorithm similar to what developed for rf X-band tests [194] could be developed and adapted to pulsed dc tests. Moreover, the system has been designed with viewports on the sides, which allowed the installation of CCD cameras to detect the light emitted by breakdowns, and to reconstruct their position through triangulation, as shown in Fig. 9. The location of breakdowns can then be easily correlated with their counterparts acquired through microscopic imaging of the electrodes surfaces, as illustrated in Fig. 10, and with the electrical signals acquired by the pulse generator to analyze for example their size and their spatial and temporal distributions. These diagnostics are complemented by the usual vacuum pressure recordings discussed for the anode tip original test system.
@@ III.2 Signals Produced During Breakdown
In this subsection we present data that illustrate features that are characteristic of breakdowns taken individually. These features include how applied fields collapse, the contemporaneous burst-like signals of electron, radiation and light emission, as well as vacuum spikes. These features are important for understanding how breakdowns evolve, the bottom three boxes in Fig. 1 and the subject of sections VI, VII and IX. The data shown are by necessity selected examples since most of the breakdown features are random and noisy. The examples are chosen to illustrate features clearly but with care so as not to bias the discussion. Some categories of randomness in the features has been identified to be important and these will be subsequently covered in the statistics subsection III.3.
@@
From the length of the drop-time in transmission and rise-time in reflection, as shown in Figs. 4 and 8, one can determine that the development time of a breakdown short circuit is very short, less than 10 ns from all available data. The time is given as an upper limit since the measurement is bandwidth limited by the structure, detection electronics and the signal acquisition.
@@
The electron current emitted during breakdown, the one causing the field collapse described above, can also be observed directly in rf accelerating structure experiments by Faraday cups installed along the beam axis. Most of the breakdown current is accelerated and then lost on the cavity walls, but a fraction is captured by the accelerating rf and is transported along the central beam axis, the beam aperture, out of the structure. It can then be directly measured using on-axis Faraday cups and connected to an oscilloscope. Example of such Faraday cup signals on normal pulses and pulses without breakdown are shown in [140].